OpenStudy (anonymous):

What is true about the complex number -5 + 5i? Check all that apply. A. The modulus is 10. B. It lies in quadrant 2. C. It lies in quadrant 4. D. The modulus is 5 square root of 2.

OpenStudy (shamim):

b

OpenStudy (anonymous):

any ideas for what quadrant it is in? graph it and you will see

OpenStudy (anonymous):

B is not the only correct answer...

OpenStudy (anonymous):

its B i graphed it.

OpenStudy (anonymous):

but im not sure if the other answers are correct like you said @satellite73

OpenStudy (anonymous):

ok good now do you know how to take the modulus

OpenStudy (anonymous):

i dont know how to take the modulus sorry

OpenStudy (anonymous):

\[|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

a=-5 b=5?

OpenStudy (anonymous):

yes, but you can ignore the minus sign if you like

OpenStudy (anonymous):

since it will go away when you square it

OpenStudy (anonymous):

so a and b are both 5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well actually \(a=-5\) as you said, but \((-5)^2=25\) just like \(5^2=25\)

OpenStudy (anonymous):

oh i see it lol

OpenStudy (anonymous):